# Inverse Constitutional AI: Compressing Preferences into Principles

### v1

@@ Abstract
Feedback data plays an important role in fine-tuning and evaluating state-of-the-art AI models. Often pairwise text preferences are used: given two texts, human (or AI) annotators select the “better” one. Such feedback data is widely used to align models to human preferences (e.g., reinforcement learning from human feedback), or to rank models according to human preferences (e.g., Chatbot Arena). Despite its wide-spread use, prior work has demonstrated that human-annotated pairwise text preference data often exhibits unintended biases. For example, human annotators have been shown to prefer assertive over truthful texts in certain contexts. Models trained or evaluated on this data may implicitly encode these biases in a manner hard to identify. In this paper, we formulate the interpretation of existing pairwise text preference data as a compression task: the Inverse Constitutional AI (ICAI) problem. In constitutional AI, a set of principles (or constitution) is used to provide feedback and fine-tune AI models. The ICAI problem inverts this process: given a dataset of feedback, we aim to extract a constitution that best enables a large language model (LLM) to reconstruct the original annotations. We propose a corresponding initial ICAI algorithm and validate its generated constitutions quantitatively based on reconstructed annotations. Generated constitutions have many potential use-cases—they may help identify undesirable biases, scale feedback to unseen data or assist with adapting LLMs to individual user preferences. We demonstrate our approach on a variety of datasets: (a) synthetic feedback datasets with known underlying principles; (b) the AlpacaEval dataset of cross-annotated human feedback; and (c) the crowdsourced Chatbot Arena data set. We release the code for our algorithm and experiments at https://github.com/rdnfn/icai.
Figure 1: The Inverse Constitutional AI problem. Starting from a set of pairwise preference feedback, we derive a set of natural language principles (a constitution) that explain the preference data. For validation, we re-construct the original preferences with an LLM judging according to the generated constitution. The constitution represents a (highly compact) compression of the preferences.
@@ 1 Introduction
State-of-the-art large language models (LLMs) rely heavily on human feedback for training and evaluation. This feedback, often in the form of pairwise text preferences, is crucial to assess advanced capabilities hard to evaluate automatically. Strategies for training on such data have seen widespread adoption, with notable examples including reinforcement learning from human feedback (RLHF) [Ouyang et al., 2022, Stiennon et al., 2020] and direct preference optimization (DPO) [Rafailov et al., 2023]. Beyond training, pairwise text preferences are also used for evaluating LLMs, such as in the Chatbot Arena [Chiang et al., 2024], where crowdsourced preferences determine rankings. Here, users interact with two anonymous models simultaneously via a web interface and select the preferred output. The resulting ranking may offer a better estimation of real-world human preferences than conventional static benchmarks, since it captures the multi-faceted nature of human preferences [Xu et al., 2023].
Unfortunately, human feedback is not without its flaws. Systematic biases in human judgement have been documented extensively in the psychology literature [Tversky and Kahneman, 1974]. It is therefore unsurprising that the human feedback used to guide and evaluate LLMs exhibits biases as well [Hosking et al., 2024, Wu and Aji, 2023, Bansal et al., 2024, Sharma et al., 2023, Xu et al., 2023]. For example, Hosking et al. [2024] find that human annotators can sometimes favour assertiveness over truthfulness. Wu and Aji [2023] observe that annotators sometimes place more importance on grammatical errors than truthfulness. Feedback data with unintended biases can be problematic: when used for fine-tuning, biased data may lead to models that exhibit the same biases. Similarly, leaderboards based on biased data will favour misaligned models [Dubois et al., 2023, 2024]. As such, it is valuable to understand the implicit rules and biases guiding annotators of feedback data. To date, however, few tools exist to detect biases in pre-existing preference data at scale. Prior work usually builds on specially designed datasets to observe biases and cannot be directly applied to pre-existing data or data generated in less controlled settings.
In this paper, we propose a novel approach to understanding preference corpora: Inverse Constitutional AI (ICAI). We make the following contributions:
@@
    Formulating the Inverse Constitutional AI (ICAI) problem. In Constitutional AI [Bai et al., 2022], a set of principles (or constitution) is used to provide feedback and fine-tune language models. The ICAI problem inverts this process: given a dataset of feedback by a human or model, we seek to compress the annotations into a set of principles that enable reconstruction of the annotations (Fig. 1).
@@
    Proposing an initial ICAI algorithm. We introduce a first ICAI algorithm that generates a set of principles based on a feedback dataset. We validate the constitutions generated by our algorithm based on their ability to help reconstruct feedback. Given the complexity of human judgement, the constitution necessarily represents a “lossy”, non-unique compression of the feedback data. Nevertheless, the interpretable nature of the principles may enable a number of promising downstream use-cases: (a) highlighting potential issues in preference data; (b) creating interpretable reward models; (c) scaling human-annotated evaluation to new models and use-cases; and (d) generating personal constitutions for customized model behavior.
@@
    Providing extensive experimental results. We test our approach experimentally on three different datasets: (a) we first provide a proof-of-concept on synthetic data with known underlying principles; (b) we then demonstrate applicability to human-annotated data on the AlpacaEval dataset [Dubois et al., 2023]; and (c) finally we showcase applicability to individual users via crowdsourced Chatbot Arena Conversations data [Zheng et al., 2023]. On the latter dataset we investigate individualized preferences, an area attracting growing interest [Kirk et al., 2024]. We demonstrate highly sample-efficient generation of individual constitutions with human-readable principles that can be edited and adapted. We release the code to reproduce our results publicly. [^1] https://github.com/rdnfn/icai
@@ 2 The Inverse Constitutional AI Problem
We describe the Inverse Constitutional AI (ICAI) problem as follows: given a set of pairwise preference feedback, the goal is to generate a corresponding constitution of natural-language principles that enable an LLM annotator to reconstruct the original preferences as well as possible. More formally, we consider the following optimization problem:
$\operatorname*{argmax}_{\text{c}}\,\{\,\text{agreement}(p_{o},p(c))\text{ s.t.}|c|\leq n\}\,,$ (1)
where $p_{o}$ are the original preferences and $p(c)$ are constitutional preferences over a pairwise text corpus $T$, generated by LLM $M$ using the constitution, $c$. The constitution is subject to the constraint of consisting of up to $n$ natural language principles that are human-readable. Agreement is defined as the percentage of constitutional preferences $p(c)$ that are identical to the original preferences $p_{o}$ (i.e. select the same output in a pair from corpus $T$). A constitution with high agreement can help interpret a preference dataset to gain insight into the underlying biases and preferences of human or AI annotators. The constitution may also be used for future preference synthesis, with an interpretable and editable set of principles.
@@ 3 Method
Figure 2: Overview of our Inverse Constitutional AI (ICAI) algorithm. Given a dataset of pairwise rankings, in Step 1 candidate principles are generated using an LLM. In Step 2, these principles are clustered using an embedding model. In Step 3, similar principles are “de-duplicated” by sampling one principle per cluster. In Step 4, each principle is tested to evaluate its ability to help an LLM reconstruct the original annotations. Finally in Step 5, the principles are filtered according to the testing results and a set of filtered principles are returned as the final constitution. Optionally, this last step is augmented with additional clustering and subsampling steps to ensure diverse principles.
Our proposed first Inverse Constitutional AI (ICAI) algorithm, outlined in Fig. 2, consists of five main steps: principle generation, principle clustering, principle sub-sampling, principle testing, and principle filtering. In the following, we will describe each step in detail.
@@
We extract candidate principles using an LLM with access to the feedback data. The principles are generated on a per-comparison basis: an LLM is prompted with a pair of texts and corresponding preference, and then asked to propose principles that explain the preference (see Section B.1 for prompts). The generated principles are in the form of natural language instructions that annotators can use to make preference decisions (e.g., “select the more polite output”). We generate a large number of candidate principles to ensure that we cover a wide range of potential rules. To ensure principle diversity, we further use multiple passes (by default 2) over the preferences with different generation prompts. The generation prompts affect the type of principles that get generated and tested (e.g., whether specific/general, positive/negative rules). Depending on the use-case, different generation prompts may be helpful.
@@
Since the first step generates a large number of candidate principles independently, almost identical principles may be generated multiple times. We then use K-means-based clustering to identify principles that are similar and can be merged into a single principle. The parameter $k$ determines the number principles considered downstream and thus affects overall computational cost.
Step 3: Principle sub-sampling.
In the third step, we de-duplicate the principles by randomly sampling one principle per cluster, leading to a diverse set of remaining principles.
@@
The fourth step evaluates the generated principles’ ability to help an LLM reconstruct the original annotations. We prompt the LLM with the generated principles to determine their support for each pair (see Section B.2). We parallelize this step, prompting the LLM with a pair of texts and multiple principles to provide a separate response (first preferred, second preferred, not relevant) for each principle. This parallelization reduces the tokens requirements compared spearate testing. While LLMs can exhibit anchoring effects when predicting multiple labels in one output [Stureborg et al., 2024], we hypothesize this effect is less pronounced for relative preferences and our experimental results indicate sufficient reliability on our datasets. We compare these results to the original labels and count the correct, incorrect, and not relevant labels for each principle separately, thereby identifying principles that help the LLM to correctly annotate the dataset.
@@
Finally, the principles are filtered based on the results of the previous testing step. We only keep principles that improve the reconstruction loss, while discarding principles that do not help or even hinder the reconstruction. We further discard principles that are marked as relevant on less than $x\%$ of the data (default $10\%$), to avoid overly specific principles that do not generalize. We order the principles according to their net contribution to correctly annotating the dataset (# correct minus incorrect annotations). We then select the top $n$ principles [^2] Experimental results for varying $n$ values available in Appendix A. The parameter $n$ may be seen as determining regularisation in our algorithm: a small $n$ may be considered highly regularised, limiting the amount of overfitting to the data possible. A large $n$ enables including more fine-grained principles that only apply to smaller subset of examples. Note that, depending on the use-case, overfitting to the training data is not necessarily a problem (e.g., for data interpretability). according to this order. Optionally, to increase principle diversity, we cluster the top $m$ ($>n$) principles into $n$ clusters as before, and sub-sample the highest ordered principle from each cluster. [^3] During our experiments we found that it is important not to be too restrictive with the number of clusters in Step 2, as good principles may never be tested. By instead filtering again for duplicates after testing, we can increase the probability of the best candidate principle passing through our algorithm (at a slightly higher testing token cost). The final ordered list of principles from this filtering step are returned as the constitution.
Inference. Given a constitution, we can validate its ability to “explain” the original feedback dataset. We do this validation using AI annotators prompted with the constitution, an approach pioneered by Bai et al. [2022] and commonly referred to as constitutional AI. Notably this leaves room for interpretation of the constitution by the AI annotator, as the constitution may be ambiguous, contradictory or incomplete. It is also dependent on the exact phrasing of the prompt and the constitution, an effect extensively studie by Li et al. [2024b] on whose work we build. When seen as an explainability tool, the possibility of conducting inference with the generated constitution is a key benefit of our approach, as it allows for validation of the explanation on a test set. This constitutional AI annotator-based inference enables us to quantitatively test the validity of our generated constitutions and their ability to explain the data while also enabling downstream use-cases such as personalized preference models.
@@ 4 Experiments
We conduct experiments on three datasets: (1) synthetic data to demonstrate the basic functionality of our algorithm, (2) human-annotated AlpacaEval data to demonstrate the applicability of our algorithm to real-world data, and (3) Chatbot Arena data to illustrate the application of our algorithm to infer individual user preferences. Throughout our experiments we primarily use two models from OpenAI: gpt-3.5-turbo-0125 (referred to as GPT-3.5-Turbo) and gpt-4o-2024-05-13 (referred to as GPT-4o). [^4] See this page for a detailed model description: https://platform.openai.com/docs/models/ We include the full numerical results in Appendix D.
Annotators. We use the AlpacaEval [^5] https://github.com/tatsu-lab/alpaca_eval, license: Apache-2.0 [Li et al., 2024b] package and their annotators as the baselines (see Section B.4). These annotators have been shown to strongly correlate with human preferences. To evaluate constitution effectiveness, we create further custom prompts that ask the model to annotate according to the principles in the constitution (see prompts in Section B.3). All our annotator outputs are parsed using functional APIs (where supported), as prior AlpacaEval experiments show this improves annotator performance. We show random baseline performance as a grey dashed line at $50\%$ in all plots.
4.1 Synthetic data as a proof of concept
Figure 3: Results on Synthetic Data. Using limited data and without fine-tuning, our constitutional annotators can reconstruct various kinds of preferences. We demonstrate our algorithm’s adaptability on three synthetic datasets (30 preferences each): one orthogonal to the base LLM’s learned preferences, one aligned with those preferences and one unaligned with them, flipping the aligned dataset. We generate constitutions for each and report agreement with the original annotations of a default LLM annotator (prompted to select the “best” output) and a constitutional annotator (prompted with a constitution). Our constitutions notably improve agreement in the orthogonal and unaligned cases and retain high agreement in the aligned case, albeit with more variance. Our method’s ability to detect biases is illustrated by the example constitution in the unaligned case. Examples of generated constitutions were chosen for illustrative purposes. Experiments use GPT-3.5-Turbo, values and error bars report mean and standard deviation over six random seeds.
We first apply our algorithm to three synthetic datasets created according to known rules intended to be aligned, unaligned, and orthogonal to the base LLM’s preferences. We provide an overview of the dataset here, with further details in Appendix F.

### v2

@@ Abstract
Feedback data is widely used for fine-tuning and evaluating state-of-the-art AI models. Pairwise text preferences, where human or AI annotators select the “better” of two options, are particularly common. Such preferences are used to train (reward) models or to rank models with aggregate statistics. For many applications it is desirable to understand annotator preferences in addition to modelling them — not least because extensive prior work has shown various unintended biases in preference datasets. Yet, preference datasets remain challenging to interpret. Neither black-box reward models nor statistics can answer why one text is preferred over another. Manual interpretation of the numerous (long) response pairs is usually equally infeasible. In this paper, we introduce the Inverse Constitutional AI (ICAI) problem, formulating the interpretation of pairwise text preference data as a compression task. In constitutional AI, a set of principles (a constitution) is used to provide feedback and fine-tune AI models. ICAI inverts this process: given a feedback dataset, we aim to extract a constitution that best enables a large language model (LLM) to reconstruct the original annotations. We propose a corresponding ICAI algorithm and validate its generated constitutions quantitatively based on annotation reconstruction accuracy on several datasets: (a) synthetic feedback data with known principles; (b) AlpacaEval cross-annotated human feedback data; (c) crowdsourced Chatbot Arena data; and (d) PRISM data from diverse demographic groups. As an example application, we further demonstrate the detection of biases in human feedback data. As a short and interpretable representation of the original dataset, generated constitutions have many potential use cases: they may help identify undesirable annotator biases, better understand model performance, scale feedback to unseen data, or assist with adapting AI models to individual user or group preferences. We release the source code for our algorithm and experiments at https://github.com/rdnfn/icai.
Figure 1: The Inverse Constitutional AI problem. Starting with pairwise preference feedback data, we derive a set of natural language principles (a constitution) that explain the preferences. For validation, we reconstruct the original preferences with an LLM judging according to the generated constitution. The constitution represents a (highly compact) compression of the preferences.
@@ 1 Introduction
State-of-the-art AI models, in particular large language models (LLMs), rely heavily on human feedback for training and evaluation. This feedback, often in the form of pairwise text preferences, is crucial to assess advanced capabilities, which are hard to evaluate automatically. Pairwise text preferences typically consist of a prompt, two model responses and an annotation selecting the “better” response. Strategies for training on such data have seen widespread adoption, with notable examples including reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020) and direct preference optimization (DPO) (Rafailov et al., 2023). Beyond training, pairwise text preferences are also widely-used for evaluating LLMs, such as in the Chatbot Arena (Chiang et al., 2024), where crowdsourced preferences determine rankings — possibly reflecting the complexity of human preferences better than conventional benchmarks (Xu et al., 2023).
However, interpreting such pairwise data is challenging: describing what exactly we train a model to do when applying RLHF with a large number of preference pairs is hard. Similarly, understanding why a model is ranked higher in a pairwise data-based leaderboard remains difficult. Yet, understanding such data is critical: human feedback is not without its flaws. Systematic biases in human judgement have been documented extensively in the psychology literature (Tversky & Kahneman, 1974). Perhaps unsurprisingly, biases have been similarly observed in the human feedback used to guide and evaluate LLMs. For example, human annotators have been observed to sometimes favour assertiveness (Hosking et al., 2024) or grammatical correctness (Wu & Aji, 2023) over truthfulness.
Feedback data with unintended biases can be problematic: when used for fine-tuning, biased data may lead to models that exhibit the same biases. Similarly, leaderboards based on biased data will favour misaligned models (Dubois et al., 2023; 2024). As such, understanding the implicit rules and biases guiding annotators of feedback data is valuable. To date, however, few general tools exist to detect biases in pre-existing preference data at scale. Prior work detecting biases usually either requires specially collected data or focuses on biases easy to programmatically measure (e.g. length bias (Dubois et al., 2024)) — difficult to extend to pre-existing datasets or less controlled settings.
In this paper, we propose a novel general approach to understanding preference corpora: Inverse Constitutional AI (ICAI). We make the following contributions:
@@
    Formulating the Inverse Constitutional AI (ICAI) problem. In Constitutional AI (Bai et al., 2022b), a set of principles (or constitution) is used to provide feedback and fine-tune language models. ICAI inverts this process: given a dataset of human or AI feedback, we seek to compress the annotations into a set of principles that enable reconstruction of the annotations (Figure 1).
@@
    Proposing an initial ICAI algorithm. We introduce a first ICAI algorithm that generates a set of principles based on a feedback dataset. We validate the constitutions generated by our algorithm based on their ability to help reconstruct feedback. Given the complexity of human judgement, the constitution necessarily represents a “lossy”, non-unique compression of the feedback data. Nevertheless, the interpretable nature of the principles may enable a number of promising downstream use cases: (a) highlighting potential issues in preference data; (b) creating interpretable reward models; (c) scaling human-annotated evaluation to new models and use cases; and (d) generating personal constitutions for customized model behaviour.
@@
    Providing experimental results and case studies. We test our approach experimentally on four datasets: (a) we first provide a proof-of-concept on synthetic data with known underlying principles; (b) we then demonstrate applicability to human-annotated data on the AlpacaEval dataset (Dubois et al., 2023); (c) we showcase applicability to interpreting individual user preferences via Chatbot Arena Conversations data (Zheng et al., 2023); (d) we investigate the use-case of bias detection on different datasets; and finally (e) we demonstrate our method’s ability to help interpret differing group preferences on PRISM data (Kirk et al., 2024). We demonstrate the highly sample-efficient generation of personalised constitutions with human-readable and editable principles. We release our code at https://github.com/rdnfn/icai.
@@ 2 The Inverse Constitutional AI Problem
Given a set of pairwise preference feedback, the Inverse Constitutional AI (ICAI) problem is to generate a corresponding constitution of natural-language principles that enable an LLM annotator to reconstruct the original preferences as well as possible. Formally, we seek to find
$\operatorname*{arg\,max}_{\text{c}}\,\{\,\text{agreement}(p_{o},p_{M}(c))\text{ s.t. }|c|\leq n\}\,,$ (1)
where $p_{o}$ are the original preferences and $p_{M}(c)$ are constitutional preferences over a pairwise preference corpus $T$, generated by LLM $M$ using the constitution $c$. The constitution is constrainted to consist of up to $n$ human-readable natural language principles. Agreement is defined as the percentage of constitutional preferences $p_{M}(c)$ identical to the original preferences $p_{o}$. A constitution with high agreement can help interpret a preference dataset to gain insight into the underlying annotator preferences and biases. The constitution may also be used for future preference synthesis, with an interpretable and editable set of principles.
@@ 3 Method
Figure 2: Overview of our Inverse Constitutional AI (ICAI) algorithm. Given a dataset of pairwise comparisons, in Step 1 candidate principles are generated using an LLM. In Step 2, these principles are clustered using an embedding model. In Step 3, similar principles are deduplicated by sampling one principle per cluster. In Step 4, each principle is tested to evaluate its ability to help an LLM reconstruct the original annotations. Finally, in Step 5, the principles are filtered according to the testing results, and a set of filtered principles are returned as the final constitution. Optionally, a final step of additional clustering and subsampling can follow to ensure diverse principles.
We propose a first Inverse Constitutional AI (ICAI) algorithm, outlined in Figure 2, consisting of five main steps: principle generation, principle clustering, principle subsampling, principle testing, and principle filtering. In the following, we describe each step in detail.
@@
We extract candidate principles using an LLM with access to the feedback data. The principles are generated on a per-comparison basis: an LLM is prompted with a pair of texts and corresponding preference, and then asked to propose principles that explain the preference (prompts in Section G.1). The generated principles are in the form of natural language instructions that inform preference decisions (e.g., “select the more polite output”). We generate a large number of candidate principles using multiple (by default 2) generation prompts and multiple principles per prompt to cover a wide range of potential rules. The generation prompts affect the type of principles that get generated and tested (e.g., specific/general, positive/negative rules).
@@
Since the first step generates a large number of candidate principles independently, almost identical principles may be generated multiple times. We use $k$-means-based clustering on embeddings to identify principles that are similar for merging. The parameter $k$ determines the number principles considered downstream and thus affects overall computational cost.
Step 3: Principle subsampling.
In the third step, we deduplicate the principles by randomly sampling one principle per cluster, leading to a diverse set of remaining principles.
@@
The fourth step evaluates the generated principles’ ability to help an LLM reconstruct the original annotations. We prompt the LLM with the generated principles to determine the ‘votes’ each principle casts across comparisons, which we then compare to the true annotations (see Section G.2). We parallelize this step, prompting the LLM with a pair of texts and multiple principles to provide a separate response (first preferred, second preferred, not relevant) for each principle. This parallelization reduces the token requirements compared to separate testing. While LLMs can exhibit anchoring effects when predicting multiple labels in one output (Stureborg et al., 2024), we hypothesize this effect is less pronounced for relative preferences and our experimental results indicate sufficient reliability on our datasets. We compare these results to the original labels and count the correct, incorrect, and not relevant labels for each principle separately, thereby identifying principles that help the LLM to correctly annotate the dataset.
@@
Finally, the principles are filtered based on the results of the previous testing step. We only keep principles that improve the reconstruction loss, while discarding principles that do not help or even hinder the reconstruction. We further discard principles that are marked as relevant on less than $x\%$ of the data (default $10\%$), to avoid overly specific principles that do not generalize. We order the principles according to their net contribution to correctly annotating the dataset (#$\text{\#correct}-\text{\#incorrect}#$ annotations). We then select the top $n$ principles according to this order (see Section F.5 for further discussion). Optionally, to increase principle diversity, we cluster the top $m$ ($>n$) principles into $n$ clusters as before, and subsample the highest ordered principle from each cluster. [^1] We found it important not to be too restrictive with the number of clusters in Step 2, as good principles may never be tested. More clusters can lead to duplicates in the tested rules, necessitating another filtering step. The final ordered list of principlesis returned as the constitution.
Inference. Given a constitution, we can validate its ability to “explain” the original feedback dataset. We do this validation using AI annotators prompted with the constitution, an approach pioneered by Bai et al. (2022b) and commonly referred to as constitutional AI. Notably this method leaves room for interpretation of the constitution by the AI annotator, as the constitution may be ambiguous, contradictory or incomplete, but results may also vary depending on the exact prompt and constitution. We build on the AlpacaEval framework (Li et al., 2024c) for our constitutional annotators. This inference using constitutional AI annotators enables quantitatively testing the validity of our generated constitutions and their ability to explain the data while also enabling downstream use cases, such as personalized preference models.
@@ 4 Experiments
We conduct experiments on four datasets: (1) synthetic data to demonstrate the basic functionality of our algorithm, (2) human-annotated AlpacaEval data to demonstrate the applicability of our algorithm to real-world data, (3) Chatbot Arena data to illustrate the application of our algorithm to infer individual user preferences, and (4) PRISM data to showcase interpreting group preferences with our algorithm. Full dataset details are available in Appendix A. We primarily use two models from OpenAI: GPT-3.5-Turbo and GPT-4o. Example constitutions in all figures were chosen for illustrative purposes. We provide more constitutions, experiment details (including numerical results), and model details in Appendices H, F and D, respectively.
Annotators. We use annotators from the AlpacaEval framework (Li et al., 2024c) as baselines that have been shown to strongly correlate with human preferences (referred to as Default annotators, see Sections G.4 and F.8). To evaluate constitution effectiveness, we create custom prompts that ask the model to annotate according to principles in a constitution (see Section G.3). The Default annotators cannot adjust to different datasets, performing poorly when preferences deviate from its default preferences. In contrast, our constitutional annotators are able to adapt, a key advantage of our approach. We show random baseline performance as a grey dashed line at $50\%$ in all plots. To contextualize ICAI’s performance, we compare it against additional baselines: a flipped Default annotator, a (fine-tuned) reward model, and an annotator based on PopAlign (Wang et al., 2024) hypothesizing principles during inference. Details are in Section F.8, summarized here. Non-adaptive baselines (pretrained reward model, (flipped) Default) perform well on some datasets but fail to adjust to all. The fine-tuned reward model adapts partially but underperforms our constitutional annotator in our low-data scenario. A custom-trained reward model with extensive data may surpass our method but would require significant resources and lack interpretability.
4.1 Proof-of-concept: Synthetic data
We first apply our algorithm to three synthetic datasets created according to known rules crafted to be aligned, unaligned, and orthogonal to the preferences internalized by the base LLM. Each dataset is generated from three principles, with 10 pairs per principle, resulting in 30 pairs per dataset. We provide an overview of these datasets here, with further details in Appendix I.